OpenStudy (anonymous):

Please Help give medal Factor: 4t^2-6ts+6t-9s

OpenStudy (ybarrap):

$$ 4t^2-6ts+6t-9s\\ =(4t^2-6ts)+(6t-9s)\\ =2t(2t-3s)+3(2t-3s)\\ =(2t-3s)(2t+3) $$

OpenStudy (anonymous):

thanks